# LISTN: Lexicon induction with socio-temporal nuance

### v1

@@ Abstract
Research on extremist online communities frequently utilizes linguistic analysis to explore group dynamics and behaviour. Existing studies often rely on outdated lexicons that do not capture the evolving nature of in-group language, nor the social structure of the community. This paper proposes a novel method for inducing in-group lexicons which incorporates its socio-temporal context. Using dynamic word and user embeddings trained on conversations from online anti-women communities, our approach outperforms prior methods for lexicon induction. We provide a new lexicon of manosphere terms, validated by human experts, which quantifies the relevance of each term to a specific sub-community. We present novel insights on in-group language which illustrate the utility of this approach.
@@ 1 Introduction
Language cues are often used to investigate extremist movements (Baele et al., 2023; Vergani and Bliuc, 2015) or to analyse the behaviour of group members (Fernandez et al., 2018). One reason for this is that groups have a tendency to develop distinctive vernaculars, also known as in-group language. According to Drake (1980), in-group language plays two important social roles within groups: first, it can obscure the discussions of the group from out-group onlookers or moderators, which is of particular importance to deviant and extremist online communities (Chancellor et al., 2016); and second, it signals cohesion and solidarity with a group by illustrating an awareness and acceptance of their norms. Since this type of informal language evolves rapidly (Stewart and Eisenstein, 2018), the use of recent lexical innovations of the group is a strong signal of group belonging (Danescu-Niculescu-Mizil et al., 2013; Zhang et al., 2017).
@@
Studies that use hand-compiled lexicons are dependent on the skill of the lexicon builders (Havaldar et al., 2024). In the case of in-group language, there are subtle signals that are challenging to identify as an out-group member (Mendelsohn et al., 2023), especially given the scale of modern text corpora. Since manual lexicon building is expensive, there is a tendency not to update lexicons frequently. For example, Rowe and Saif (2016) include a 7 year old lexicon in their study on online pro-ISIS language. These factors impact the value of the lexicon with resulting negative effect on the quality of downstream studies.
@@
Despite the highly dynamic and social nature of groups and their language, existing methods for computationally inducing in-group language tend to focus only on linguistic information. For example, Lucy and Bamman (2021) use statistical properties of word occurrences and contextualised word embeddings. Similarly, Farrell et al. (2020) and Yoder et al. (2023) use approaches based on topic models and static word embeddings.
In this work, we propose methods for lexicon induction with socio-temporal nuance (LISTN). Our approach leverages dynamic word and user embeddings, developed in a recent work by De Kock (2024), in the context of online anti-women communities collectively known as the manosphere. Our experiments indicate that LISTN-based methods outperform a range of baselines based on prior work, achieving an average precision score of 0.77 on a newly crafted test set for lexicon induction. We use existing lexicons of manosphere language and human expert validation in our evaluation, and release a lexicon of 455 new manosphere terms with scores representing their relevance to specific manosphere subcommunities.
Using this lexicon, we further find that word embeddings of in-group terms tend to be more static (i.e. showing less variation across timesteps) compared to words of similar frequencies in the general vocabulary. We further investigate the interaction of in-group language and violent language in the manosphere, finding that there is a positive correlation ($\rho=0.53$) but that the strength of the relationship varies for different subcommunities. Finally, we look at the overlap in vernacular between the different groups, finding that Incels are the most distinctive in their in-group language. Our findings provide quantitative support for prior claims in sociological studies of relationships between different sub-groups of the manosphere.
2 Socio-temporal models of language
As mentioned in Section 1, we leverage word and user embeddings from prior work. The architecture of Yao et al. (2018) deconstructs user-content representations over different timesteps, which yields dynamic user embeddings and word embeddings. A recent work by De Kock (2024) extended this model, proposing a joint model for community structure and language, which accounts for temporal evolution in both factors. They show that the resulting embeddings capture finegrained changes in the relevance of a term to a community, and to sub-groups within the community.
To induce lexicons with such a joint model, we use pretrained embeddings for both of the abovementioned architectures, as produced by De Kock (2024), in this work. The models are pretrained on data from the manosphere, which is known to have highly distinctive and dynamic in-group language (Bogetić, 2023; Farrell et al., 2020) and sub-group structure (Ribeiro et al., 2021), making it well-suited for our focus on socio-temporal context. Below we review the architectures and the training data used. Full details are provided in the respective papers.
@@ 2 Socio-temporal models of language
A user-content matrix $C_{t}$ is constructed, using pointwise mutual information (PMI) to capture how frequently a user uses a given word compared to a background corpus during timestep $t$. Matrix factorisation is used to obtain user embeddings $U_{t}$ and word embeddings $W_{t}$ such that $C_{T}\approx U_{t}\cdot W_{t}^{T}$, for all $t$. To ensure alignment of embeddings across timesteps, the learning objective includes a regularisation term that penalises large inter-timestep changes.
@@
To capture interpersonal dynamics and social structure, a user-user adjacency matrix $A_{t}$ is constructed to capture how frequently two users interact in the same thread at time $t$. This matrix is jointly decomposed with $C_{t}$ to yield $U_{T}$, $V_{T}$ and $W_{T}$ such that $C_{T}\approx U_{t}\cdot W_{t}^{T}$ and $A_{T}\approx U_{t}\cdot V_{t}^{T}$. A similar smoothing mechanism is used to align embeddings over time. De Kock (2024) show that the joint factorisation outperforms the content-only approach in their evaluation.
Data
The models are trained on over 4 million utterances from 50 manosphere subreddits over 9 months, using the Reddit portion of the corpus of Ribeiro et al. (2021). The subreddits in the dataset are mapped to five different manosphere sub-communities, referred to as categories in this paper: Incels, Men’s Rights Activists (MRA), Men Going Their Own Way MGTOW, Pick-up Artists (PuA), and The Red Pill (TrP). The dataset further includes data from two anti-manosphere communities (referred to as Criticism) and three Mental Health related communities, which we use as a control group in this study.
Embeddings are produced at monthly timesteps, representing April-December 2018. The data is filtered to include only users who are active for 3 or more timesteps and words that are used by more than 20 users. The PMI in the $C$ matrix is calculated against a background corpus of 40 million Reddit posts by Dziri et al. (2019) is used. For tokenisation, the NLTK word tokeniser is used. Words are lowercased and non-alphabetic characters are removed.
@@ 3 LISTN
Using the models described in Section 2, we evaluate a number of approaches for inducing lexicons. We refer to these methods as LISTN: lexicon-induction with socio-temporal nuance, and distinguish between LISTN-CA (jointly training on content and adjacency matrices) and LISTN-C (using only the content matrix).
Recall that the user-content matrix $C$ is constructed using PMI to compare how often a user uses a term compared to a large background corpus. To use the resulting word embeddings $w$ and user embeddings $u$ for lexicon induction, we evaluate several methods based on the low-rank reconstruction $u_{t}\cdot w_{t}^{T}$, referred to as the relevance of a word to a user by De Kock (2024). The rationale is that we hope to find words which are used more often than expected (as indicated by PMI), using embeddings to provide a more nuanced view of a word or user compared to word frequencies. Farrell et al. (2020) find that the usage of in-group terms varies between the different manosphere subcommunities.
Therefore, we explore several methods that allow for sub-group specialisation.
@@
We calculate a vector of word relevance scores $y_{t}=\overline{u}_{t}\cdot w_{t}^{T}$, where $\overline{u}_{t}$ represents the centroid of all users in the training set at time $t$.
@@
We calculate the centroid of each of the manosphere categories (as described in Section 3) $\overline{c}_{t}$ as the mean of all users who contributed to that category at timestep $t$. For every word, we then find its maximum relevance score across all categories: $y_{t}=\max_{c_{t}}(\overline{c}_{t}\cdot w_{t}^{T})$. We use maximum aggregation (as opposed to mean or median) because a word only needs to be highly relevant to one sub-group to be considered as in-group language.
@@
Similar to the category-based method, we find the maximum relevance score at the subreddit level using its centroid $\overline{s}_{t}$: $y_{t}=\max_{s_{t}}(\overline{s}_{t}\cdot w_{t}^{T})$.
@@
We use K-means clustering of the user embeddings to find subcommunities. The motivation is that there may be groups that span different categories or subreddits, or sub-groups within a given category. For each cluster $r$ we find $y_{t}=\max_{r_{t}}(\overline{r}_{t}\cdot w_{t}^{T})$. We do this for $K\in\{5,20,100\}$. Since the clustering dependends on random initialisation, we report the mean over 5 runs.
@@
Departing from the reconstruction-based approaches, we use a list of confirmed terms from manosphere lexicons (more details in Section 4.3) and find their nearest neighbours based on cosine similarity. The score for each neighbouring word is given by its maximum similarity to a lexicon word.
@@
These approaches produce a score for each term at every timestep. To aggregate, we use the maximum score of a term across all timesteps. This has the benefit of potentially capturing words that move into and out of popularity over the training window.
4 Experimental setup
4.1 Task definition
The goal of this work is to induce lexicons of manosphere language. For the purpose of evaluation, we focus on single-token lexical innovations. Our rationale for this task definition is provided below.
A challenge in evaluating lexicons is that they are not a precise science and often rely on subjective expert judgment, as discussed in Section 1. For example, the lexica in Table 3 contain words such as hurt and dumb, which represent borderline cases in deciding whether they are sufficiently relevant for inclusion. A further challenge is that dogwhistles (words that take on a special meaning within an in-group that is different from its meaning to the general population) are difficult for researchers to identify (Lucy and Bamman, 2021; Mendelsohn et al., 2023).
To address these challenges, we focus on lexical innovations in our evaluation. Lexical innovations are a form of in-group language which involve non-standard or extragrammatical words (Slotta, 2016); for example, “foid” is a common derogatory term for a woman in the Incel community. Lexical innovations provide a clearer decision boundary as they tend to be more easily attributed to a specific community. Notably, this is a constraint on the evaluation and not a limitation of the method; all methods used in this work can produce scores for standard and non-standard language alike.
A set of standard words are required in order to identify non-standard words, or lexical innovations. For this purpose, we use the vocabulary of GloVe-400k (Pennington et al., 2014), which is based on news and Wikipedia articles. We make an exception to include in-vocabulary words from existing manosphere lexicons in our evaluation, irrespective of whether they are lexical innovations. Since prior work has deemed these terms to be of particular importance to the manosphere, we consider them to be relevant and useful for evaluation.
Finally, we consider only single-token words to limit the vocabulary size. Future work may investigate N-gram models, since multi-word in-group expressions are also observed in the data. However, the common multi-word expressions are often made into acronyms; for example, AWALT is an acronym of “all women are like that”.
@@ 4.2 Baselines
Embedding expansion
We train a word2vec model (Mikolov et al., 2013) on the same subset of posts used to train the LISTN models. We then use a bootstrap method to find the nearest neighbours of a lexicon of known manosphere terms using cosine similarity. Similar approaches are followed by Yoder et al. (2023), Farrell et al. (2020) and Havaldar et al. (2024) for lexicon expansion.
@@
Similar to Lucy and Bamman (2021) and Zhang et al. (2017), we evaluate variants of PMI, including normalised PMI (NPMI). Given a word $w$ and a particular setting (e.g. subreddit or community) $s$:
$\text{PMI}(w,s)=\text{log}\frac{P(w|s)}{P(w)},$ (1)
$\text{NPMI}(w,s)=\frac{\text{PMI}(w,s)}{-\text{log}P(w,s)}.$ (2)
Probabilities are calculated based on word frequencies. To determine $P(w)$ and $P(w,s)$, we use the same background corpus used to construct the $C$ matrix. The (N)PMI of a term is then calculated based on its probability in our training set compared to the background corpus. To account for potential sub-group specialisation, we also evaluate (N)PMI on a subreddit and category level. We use the maximum PMI for a given word across the set of subcommunities to produce its score. Finally, we evaluate the (N)PMI of the full community using monthly slices of data and maximum aggregation.
4.3 Existing lexicons
We use existing lexicons of manosphere language as basis for evaluating our methods. The lexicons and the number of terms, filtered to include only terms that are in our corpus, are shown in Table 1.

### v3

# Inducing lexicons of in-group language with socio-temporal context

@@ Abstract
In-group language is an important signifier of group dynamics. This paper proposes a novel method for inducing lexicons of in-group language, which incorporates its socio-temporal context. Existing methods for lexicon induction do not capture the evolving nature of in-group language, nor the social structure of the community. Using dynamic word and user embeddings trained on conversations from online anti-women communities, our approach outperforms prior methods for lexicon induction. We develop a test set for the task of lexicon induction and a new lexicon of manosphere language, validated by human experts, which quantifies the relevance of each term to a specific sub-community at a given point in time. Finally, we present novel insights on in-group language which illustrate the utility of this approach.
Inducing lexicons of in-group language
with socio-temporal context
Christine de Kock University of Melbourne Computing and Information Systems christine.dekock@unimelb.edu.au
@@ 1 Introduction
Social groups have a tendency to develop distinctive vernaculars, also known as in-group language. According to Drake (1980), in-group language plays two important social roles within groups: first, it can obscure the discussions of the group from out-group onlookers or moderators, and second, it signals cohesion and solidarity with a group by illustrating an awareness and acceptance of their norms. Since this type of informal language evolves rapidly (Stewart and Eisenstein, 2018), the use of recent lexical innovations of the group is a strong signal of group belonging (Danescu-Niculescu-Mizil et al., 2013; Zhang et al., 2017).
@@
Due to its strong relation to implicit social dynamics, lexicons of in-group language are frequently used to study communities (e.g. Baele et al., 2023; Rowe and Saif, 2016). Studies that use hand-compiled lexicons are dependent on the skill of the lexicon builders (Havaldar et al., 2024). In the case of in-group language, there are subtle signals that are challenging to identify as an out-group member (Mendelsohn et al., 2023), especially given the scale of modern text corpora. Since manual lexicon building is expensive, there is a tendency not to update lexicons frequently. For example, Rowe and Saif (2016) include a 7-year-old lexicon in their study on online pro-ISIS language. These factors impact the quality of lexicons, with resulting negative effect on the quality of downstream studies.
@@
Prior work in NLP has proposed computational methods for inducing lexicons of in-group language, often framing the task as identifying relevant words using large text corpora from the group in question. For example, Lucy and Bamman (2021) use statistical properties of word occurrences and contextualised word embeddings. Similarly, Farrell et al. (2020) and Yoder et al. (2023) use approaches based on topic models and static word embeddings. However, existing methods for computationally identifying in-group language tend to focus only on linguistic information, despite the highly dynamic and social nature of groups and their language.
In this work, we propose and evaluate methods for Lexicon Induction with Socio-Temporal Nuance (LISTN) using dynamic word- and user embeddings. This framework provides a natural way to integrate statistical properties of the text with the social structure of the community, by modelling users and words in the same space. Our experiments indicate that LISTN-based methods outperform a range of baselines based on prior work, achieving an average precision score of 0.77 on a newly crafted test set for lexicon induction [^1] Github: github.com/unimelb-nlp/listn.
Our study centers on the manosphere, a growing movement of anti-women online communities which has been described as a violent extremist ideology (Baele et al., 2023) and has been associated several real-world acts of terror against women (Latimore and Coyne, 2023). These communities are exceptionally linguistically innovative (Bogetić, 2023), making them a particularly appropriate application area for this work, in addition to being a pressing current concern. The manosphere is a highly fractured collective, with well-developed and continuously evolving sub-group structure (Ribeiro et al., 2021), which also highlights the need for incorporating social and temporal dimensions for lexicon induction.
To evaluate our lexicon induction methodology, we use existing lexicons of manosphere language and human expert validation to construct a novel test set. We further release a lexicon of 455 new manosphere terms with scores representing their relevance to specific manosphere subcommunities. Using this lexicon, we find that word embeddings of in-group terms tend to be more static (i.e. showing less variation over time) compared to words of similar frequencies in the general vocabulary. Finally, we look at linguistic specialisation within subgroups of the manosphere, finding that Incels are the most distinctive in their vernacular. These analyses illustrate the utility of our approach for studying in-group language as a phenomenon, and for drawing out latent group dynamics as signalled through language.
2 Related work
A number of prior works have investigated in-group language, its evolution over time, and its relation to social dynamics. Chancellor et al. (2016) investigate linguistic variation and the processes by which existing words are modified to create neologisms in the context of pro-anorexia groups, finding that lexical innovations are associated with increased engagement and more extreme content. Danescu-Niculescu-Mizil et al. (2013) explored the role of linguistic evolution in groups, using snapshot n-gram language models to measure the deviation of a user from the community. They illustrate that the non-adoption of lexical innovations is predictive of user churn. Zhang et al. (2017) characterises how distinctive and temporally dynamic online communities are in their language, finding that communities with more dynamic and distinctive language are likely to have higher retention of members. Stewart and Eisenstein (2018) and Del Tredici and Fernández (2018) investigate factors that influence the dissemination of new lexical innovations, finding that linguistic and social factors both play a role in this process. These studies illustrate the intimate relationship between social dynamics and language, and the dynamic nature of in-group language and groups. Our work uses a holistic approach for incorporating these different effects to induce lexicons of in-group language.
Language cues have been used to investigate extremist movements; for instance, by Baele et al. (2023) and De Kock and Hovy (2024) in the context of the manosphere, and by Rowe and Saif (2016) to study ISIS radicalisation. However, these works rely on small, manually constructed lexicons. Through this work, we hope to enable future research on this important topic by producing lexicons with a strong empirical justification, grounded in the socio-temporal context of the community being studied.
Our methods leverage dynamic, jointly-learned word and user embeddings (described in Section 3). The representation learning methodology was introduced in our earlier study, De Kock (2024), to forecast what communities a user will interact with, to characterise the level of violent language within sub-groups of the manosphere, and to measure changes in the relevance of temporal concepts (e.g. #MeToo) to the community. Here, we reimplement and optimise the methodology (described in Section 3.2 but use the same training data. We then propose and compare 6 different methods for using these embeddings to induce lexicons.
3 Representation learning
The Cerberus architecture (De Kock, 2024) performs a shared temporal matrix factorisation over content and adjacency matrices, which are user-content and user-user representations. A content-only model variation decomposes user-content matrices over time to yield dynamic word embeddings.
@@ 3 Representation learning
At each timestep $t$, a user-content matrix $C_{t}$ is constructed using positive pointwise mutual information (PPMI), which quantifies how strongly the presence of user $i$ and word $j$ are associated beyond what would be expected if they were independent (based on a background corpus):
$\text{PMI}(i,j)$ $=\text{log}\frac{P(j|i)}{P(j)},$ (1)
$\text{PPMI}(i,j)$ $=\text{max}\big{(}\text{PMI}(i,j),0\big{)}.$ (2)
$P(j)$ calculated is based on word frequencies in a background corpus of 40 million Reddit posts by Dziri et al. (2019). Matrix factorisation is used to learn user representations $U_{t}$ and word representations $W_{t}$ such that $C_{t}\approx U_{t}\cdot W_{t}^{T}$, for all $t$.
@@
To represent interpersonal dynamics, a user-user adjacency matrix $A_{t}$ is constructed to capture how frequently two users interact in the same thread at time $t$, with L1 normalisation. This matrix is jointly decomposed with $C_{t}$ to yield $U_{t}$, $V_{t}$ and $W_{t}$ such that $C_{t}\approx U_{t}\cdot W_{t}^{T}$ and $A_{T}\approx U_{t}\cdot V_{t}^{T}$, where $V_{t}$ is the context user embedding matrix.
To ensure alignment of embeddings across timesteps, the learning objective includes a regularisation term that penalises large inter-timestep changes. The joint factorisation was found to outperform the content-only approach in De Kock (2024).
3.1 Data
The models are trained on over 4 million utterances from 50 manosphere subreddits over 9 months, using the Reddit portion of the corpus of Ribeiro et al. (2021). The subreddits in the dataset are mapped to five different manosphere sub-communities, referred to as categories in this paper: Incels, Men’s Rights Activists (MRA), Men Going Their Own Way MGTOW, Pick-up Artists (PuA), and The Red Pill (TRP). The dataset further includes data from two anti-manosphere communities (referred to as Criticism) and three Mental Health related communities, which we use as a control group in this study. Embeddings are produced in monthly windows, representing April-December 2018.
The data is filtered to include only users who are active for 3 or more timesteps and words that are used by more than 20 users, yielding 33,880 users and 44,679 words. The NLTK word tokeniser is used for tokenisation. Words are lowercased and non-alphabetic characters are removed.
3.2 Implementation
The models used in De Kock (2024) were implemented in TensorFlow to match the evaluation setting of antecedent systems, and were not optimised for scaling to large datasets. We re-implement [^2] Github: github.com/unimelb-nlp/listn these models in PyTorch, using the generalised matrix factorisation formulation (He et al., 2017) which (1) avoids reading the full sparse matrices into memory, (2) enables batch-wise updates for faster convergence, and (3) allows for increased flexibility to deploy other models for representation learning.
In this formulation, a training step is performed over a batch of users for a given timestep, rather than updating only once all the data has been seen. During the forward pass, the model computes interaction vectors for the users in the batch against the full set of context users and words for that timestep. The model parameters are updated after each tuple of (user-user, user-word) batches. Hyperparameters are tuned based on the validation loss and are shown in Table 4 in Appendix A.
@@ 4 LISTN
Using the models described in Section 3, we evaluate a number of approaches for inducing lexicons. We refer to these methods as LISTN (Lexicon Induction with Socio-Temporal Nuance) and distinguish between LISTN-CA (jointly training on Content and Adjacency matrices) and LISTN-C (using only the Content matrix).
Recall that the user-content matrix $C$ is constructed using PMI to compare how often a user uses a word compared to a large background corpus. We evaluate several methods for lexicon induction that are based on the low-rank reconstruction $\hat{C}_{t}$. The rationale is that we aim to find words that are used more often than expected (as indicated by PMI), using embeddings to provide a more nuanced view of a word or user compared to word frequencies.
Given a user embedding $u_{i,t}$ and a word embedding $w_{j,t}$, we calculate the relevance of word $j$ to user $i$ at time $t$ as ${r(i,j,t)=u_{i,t}\cdot w_{j,t}^{T}}$ (as also done in De Kock, 2024). This can be generalised to groups of users by finding the group centroid as the mean of the individual user embeddings, denoted $\overline{u}_{t}$. Farrell et al. (2020) find that the usage of in-group terms varies between the different manosphere categories. Therefore, we explore several methods that allow for sub-group specialisation. Each of the below methods produce a (scalar) score $r$ representing the relevance of a word $j$ within the manosphere ecosystem at time $t$.
@@
We calculate word relevance scores as ${r_{comm}(j,t)=\overline{u}_{t}\cdot w_{j,t}^{T}}$, where $\overline{u}_{t}$ represents the centroid of all users in the training set at time $t$.
@@
We calculate a centroid $\overline{u}_{k,t}$ for each of the manosphere categories (as described in Section 3), defined as the mean embedding of all users who contributed to category $k$ at timestep $t$. For every word, we then find its maximum relevance score across all categories: ${r_{cat}(j,t)=\max_{k}(\overline{u}_{k,t}\cdot w_{j,t}^{T})}$. We use maximum aggregation (as opposed to mean or median) because a word only needs to be highly relevant to one sub-group to be considered as valid in-group language.
@@
Similar to the category-based method, we find the maximum relevance score for a subreddit $p$ using its centroid: ${r_{sub}(j,t)=\max_{p}(\overline{u}_{p,t}\cdot w_{j,t}^{T})}$.
@@
We use K-means clustering of the user embeddings to find subcommunities. The motivation is that there may be groups that span different categories or subreddits, or sub-groups within a given category. For each cluster $n$ we find ${r_{clust}(j,t)=\max_{n}(\overline{u}_{n,t}\cdot w_{j,t}^{T})}$. We do this for $K\in\{5,20,100\}$. Since the clustering depends on random initialisation, we report the mean over 5 runs.
@@
Departing from the reconstruction-based approaches, we use a list of confirmed terms from manosphere lexicons (provided in Table 1) and find their nearest neighbours based on cosine similarity of the $W$ embeddings. The score for each neighbouring word is given by its maximum similarity to a lexicon word.
@@
Aggregation
These approaches produce a score for each word at every timestep. To aggregate, we use the maximum score of a word across all timesteps. This has the benefit of potentially capturing words that move into and out of popularity over the training window.
5 Evaluation
The goal of this work is to induce lexicons of in-group language. For the purpose of evaluation, we frame this as a binary classification task over single-token lexical innovations. Our rationale for this task definition and our evaluation setting is provided below. The test set constructed for this task is described in Section 6.
A challenge in evaluating lexicons is that they are not a precise science and often rely on subjective expert judgment (Havaldar et al., 2024), as discussed in Section 1. For example, the lexicons in Table 3 contain words such as hurt and dumb, which represent borderline cases in deciding whether they are sufficiently relevant for inclusion. A further challenge is that dogwhistles (words that take on a special meaning within an in-group that is different from its meaning to the general population) are difficult for researchers to identify (Lucy and Bamman, 2021; Mendelsohn et al., 2023).
To address these challenges, we focus on lexical innovations in our evaluation. Lexical innovations are a form of in-group language which involve non-standard or extragrammatical words (Slotta, 2016); for example, foid is a common derogatory term for a woman in the Incel community. Lexical innovations provide a clearer decision boundary as they tend to be more easily attributed to a specific community. Importantly, this is a constraint on the evaluation and not a limitation of the method; all methods used in this work can produce relevance scores for standard and non-standard language alike.
A set of standard words is required in order to identify non-standard lexical innovations. For this purpose, we use the vocabulary of GloVe-400k (Pennington et al., 2014), which is based on news and Wikipedia articles. Words that are not in this vocabulary are considered non-standard. We make an exception to include in-vocabulary words from existing manosphere lexicons (detailed in Section 6) in our evaluation, irrespective of whether they are lexical innovations. Since prior work has deemed these terms to be of particular importance to the manosphere, we consider them to be relevant and useful for evaluation.
Finally, we consider only single-token words to limit the vocabulary size. Future work may investigate N-gram models, since multi-word in-group expressions are also observed in the data. Within the manosphere, common multi-word expressions are often made into acronyms; for example, AWALT is an acronym of “all women are like that”.
@@ 5.1 Baselines
Lexicon expansion
Similar to the approaches of Yoder et al. (2023), Farrell et al. (2020) and Havaldar et al. (2024), we train a word2vec model (Mikolov et al., 2013) on the same subset of posts used to train the LISTN models. We then find the nearest neighbours of known manosphere in-group language (based on the lexicons in Table 1) using cosine similarity over the $W$ embeddings.
@@
Similar to Lucy and Bamman (2021) and Zhang et al. (2017), we evaluate variants of PMI, including PPMI (Eq. 1) and normalised PMI (NPMI). Given a word $j$ and a particular setting (e.g. subreddit or community) $k$:
$\text{NPMI}(j,k)$ $=\frac{\text{PMI}(j,k)}{-\text{log}P(j,k)}.$ (3)
To determine $P(j)$ in the PMI calculation, we use the same background corpus used to construct the $C$ matrices. Probabilities are calculated based on word frequencies. Mirroring the LISTN evaluations, we account for sub-group specialisation by evaluating each PMI metric at the level of the full community, subreddit, and category, using the maximum score for a given word across the set of subgroups in each case. We also evaluate the NPMI and PPMI using monthly slices of data and maximum aggregation to mimic the temporally-aware nature of the LISTN methods.
5.2 Metrics
We report the average precision (AP) and area under the receiver-operator characteristic curve (AUROC) for each method. The AP provides a measure of the model’s precision over all possible recall values; that is, how well the model avoids false positives. AUROC is a measure of its ability to distinguish between positive and negative cases. Davis and Goadrich (2006) show that a model dominates in AUROC space if and only if it dominates in AP space, whereas the converse is not true; as such, we prioritise the AP. We use the randomised permutation test with Monte Carlo approximation ($R=9999$ and $\alpha=0.05$) for significance testing.
6 Test set construction
We perform a preliminary scoring step for the purpose of constructing a test set for our main evaluation. Existing lexicons of manosphere language are used as a reference point.
Existing lexicons
The lexicons and the number of words in each, filtered to include only words that are in our corpus, are shown in Table 1. For the Farrell et al. (2019) dataset, we follow Ribeiro et al. (2021) in excluding the lexicon categories that are not specific to the manosphere (e.g. racism). The Hatebase333hatebase.org/ and IncelWiki444incels.wiki/w/Incel_Glossary lexicons were collected by us and will be released. Combined, this provides a set of 483 unique words.